OpenStudy (anonymous):

Help ?? Write the equation in standard form. 2x+½y=3

OpenStudy (anonymous):

4x+y=6

OpenStudy (anonymous):

Thank-you.How'd you get that like step by step if you can, and is that all?

OpenStudy (anonymous):

multiply that equation by 2 we get 2*2x+2*(1/2y)=3*2 then 4x+y=6

OpenStudy (anonymous):

Ohh I see now, Thnak-you !! (:

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ohh yeaah that's right I forgot about y=mx+b, Thanks you two for the last step